# arXiv:2603.01456v1[gr-qc] 02 Mar 2026 thanks: Corresponding author

### 02 Mar 2026 v1

@@ III.1 EVP method
III.1.1 Discretization choices and convergence
@@ III.2 Leaver method + band reduction
III.2.1 Discretization choices and convergence
@@ IV Numerical Results
In this section we present QNM shifts for overtones $n=0,1,2,3$ and multipoles $\ell=2,3,4$ in both even and odd sectors. We work in units $r_{g}=1$ throughout, so that the $n=0$ modes can be compared directly with results of Ref. [1].
@@ IV.1.1 Null operators
It is instructive to compare with the results of Ref. [1] that uses method developed in [33] to compute the QNMs. They have focused on the fundamental modes ($n=0$) with $\ell=2,3,4$ and reported maximal null-operator shifts of order $10^{-4}$ in the even sector and $10^{-6}$ in the odd sector. Over the same set of modes, our EVP implementation reduces these residuals by several orders of magnitude, yielding $|\delta_{\mathrm{EVP}}|\lesssim 10^{-8}$ while simultaneously extending the null test to overtones up to $n=3$ without loss of control. We speculate that the improvement is primarily due to avoiding the series truncation error used In Ref. [1] and  [33]: in EVP we compute the shift directly from the perturbed eigenvalue problem at high precision, and verify convergence so the residual in null cases is driven down to a much lower numerical floor.
@@ IV.1.2 Control Operators
Up to an overall sign that can be traced to a difference in conventions, the EVP results for $\mathcal{O}_{9}$ and $\mathcal{O}_{10}$ operators gives a fractional shifts $\delta_{\mathrm{EVP}}$ are $\mathcal{O}(10^{0}-10^{2})$. They agree with those of Ref. [1] in magnitude for both the real and imaginary parts upon rounding to the two decimal places reported in their Table I.
@@
Finally, we stress that the EVP computation only probes the linear response of the spectrum, encoded in the first–order coefficients $\omega_{1}$ and hence in $\delta_{\mathrm{EVP}}$. By construction, this method cannot determine the range of EFT coupling $\zeta$ over which the spectrum remains linear in the sense of $\omega(\zeta)\approx\omega_{0}+\zeta\,\omega_{1}+\mathcal{O}(\zeta^{2})$ .
@@ IV.2 Leaver method + band reduction
Now compute the EFT-corrected frequency $\omega(\zeta)$ for $\zeta=\alpha\,10^{-p}$ with $p\in\{3,4,\ldots,15\}$ and $\alpha\in\left\{1,\frac{3}{4},\frac{1}{2},\frac{1}{4}\right\}$, keeping only $\zeta\in[10^{-15},10^{-3}]$.
@@
The EFT linear expansion in $\zeta$ is only meaningful as long as the background is not strongly deformed. An out-of-envelope diagnostic is the correction to the horizon radius $r_{H}$. For example, we find that at $d_{5}\zeta=10^{-2}$ the horizon is shifted by about $12\%$ from $r_{g}$, while for $d_{9}\zeta=10^{-2}$ the shift is about $10\%$. [^2] Pushing to $d_{i}\zeta=10^{-1}$ is not even physically meaningful, since $r_{H}$ would formally become negative. To remain safely within the linear regime, in what follows we therefore take $\zeta_{\max}=10^{-4}$ as our default upper limit, and treat $\zeta=10^{-3}$ only as a reference point illustrating the onset of linear perturbation failure [^3] In what follows, $d_{i}\in\{0,1\}$ is a switch that selects the $i$th EFT operator, while $\zeta$ is the explicit perturbation parameter..
@@
Although the underlying field equations are linearized in $\zeta$, the numerical procedure used to extract $\omega(\zeta)$ via Leaver’s method can effectively introduce higher-order dependence on $\zeta$. Two immediate sources are: (i) the Gaussian elimination step, which combines $\zeta$-dependent coefficients in a way that generates products and therefore higher-order terms, and (ii) the nonlinear Newton solver employed to locate the roots of the continued fraction, which can also imprint nonlinear $\zeta$-dependence through its iterative updates. As a result, the numerically extracted frequency is more accurately viewed as an expansion
@@
where $a_{1}$ is the physical linear EFT coefficient, while $a^{(1)}_{2},a^{(1)}_{3}$$\ldots$ denote effective higher-order contributions induced by the Leaver extraction procedure, even when the input perturbation is truncated at $\mathcal{O}(\zeta)$. Consequently, estimating $a_{1}$—especially in cases where the linear contribution is suppressed or vanishes for a given operator—is a subtle task.
IV.3 Diagnostics
@@
IV.3.1 What is the leading power?
@@
where the coefficient $C_{\ell n}\in\mathbb{C}$ are generally complex. On a chosen coupling window $\zeta_{\min},\zeta_{\max}]$, we then infer the leading scaling via two complementary steps: (i) estimate a continuous effective exponent $p_{\rm eff}$ from log–log slopes of the amplitude $\|\delta_{\mathrm{Leaver}}\|$ across adjacent $\zeta$ values, and we then average these local slope estimates across the $\zeta$ points within the chosen fitting window $\zeta_{\min},\zeta_{\max}]$ to obtain the reported $p_{\rm eff}$ (with scatter across the window used to define $\sigma_{p}$), and (ii) test candidate integers $p\in\{1,2,3\}$ by rescaling the data by $\zeta^{p}$ and selecting the $p_{\star}$ that makes the rescaled complex series as flat as possible. Finally, with $p_{\star}$ fixed, we obtain the corresponding complex coefficient $C$ from a least-squares fit using the flatness ratio $\sigma/\mu$ reported in the tables where $N$, $\mu$ and $\sigma$ is the sample population, mean and standard deviation respectively inside the fit window. Here is where the scaling terminology is defined. I suggest defining explicitly the term, that is, somewhere in the text one should read explicitly the terminology ”power scaling” and its definition. People woudl wonder if the power scaling of the caption is the same of the power law. An alternative is to say in the caption Power law scaling or something like this, but there should be a clear connection of terminology
@@
Upon tightening the window to $\zeta_{\max}=10^{-4}$ and $10^{-5}$, the same $p_{\star}=2$ conclusion is reinforced: $p_{\rm eff}$ moves closer to $2$ and $\sigma/\mu$ decreases, consistent with convergence toward the asymptotic quadratic regime. We therefore identify $p=2$ as the leading power for both null operators in both parities
@@
Tightening the window to $\zeta_{\max}=10^{-4}$ and $10^{-5}$ preserves the same $p_{\star}=1$ selection while reducing uncertainties and significantly decreasing $\sigma/\mu$; the fitted complex coefficient $C$ (reported via $\Re(C)\,|\,\Im(C)$) varies only mildly across windows, consistent with a well-resolved linear leading term.
IV.3.2 Polynomial fits
@@
$M_{123}$ $:\quad\Delta\omega(\zeta)\approx c^{(123)}_{1}\,\zeta+c^{(123)}_{2}\,\zeta^{2}+c^{(123)}_{3}\,\zeta^{3}.$
We use $M_{123}$ only for the null cases. Moreover, although our calculated QNMs have $\sim 60$ digits of precision and we could in principle include work with $M_{1234}$ as well, we won’t to avoid overfitting. This is the same reason why $M_{123}$ is not going to be used for control cases.
@@
IV.4 Cross Check: EVP Vs Leaver
@@ V Discussion and Conclusion
Next-generation gravitational-wave detectors will plausibly make QNMs sensitive probes of small, high-scale departures from GR in the nonlinear strong-field regime. A well-controlled way to model some beyond-GR effects is the gravitational EFT program implemented through the MTE framework. However, before trusting tiny EFT-induced deviation from QNMs as robust theoretical predictions, the framework and its numerical extraction machinery must be stress-tested: (i) the setup is intrinsically a two-parameter ($\epsilon$ and $\zeta$) perturbation problem, and (ii) the expected effects can be contaminated with numerical noise–especially for overtones–if not enough precision was used.
@@
On a Ricci-flat background, the operators $\mathcal{O}_{5}$ and $\mathcal{O}_{8}$ are removable by local field redefinitions and therefore induce no first-order ($\mathcal{O}(\zeta)$) vacuum QNM shifts in theory [1, 7]. This provides a sharp, operator-level null prediction against which both solvers can be stress-tested.
@@ Appendix B More Details On Numerical Methods
This gives a numerically stable approximation to $R_{0}(\omega)$, and we then solve the QNM condition $F(\omega)=0$ using a two-point complex secant method. The precise numerical tolerances and parameter choices for $N_{\rm tail}$, $\Delta N$, and $p$ are summarized in the LABEL:Discretization_choices_and_convergence.

### 07 Mar 2026 v2

# thanks: Corresponding author

@@ III.1 EVP method
III.1.1 EVP Discretization choices and convergence
@@ III.2 Leaver method + band reduction
III.2.1 Leaver Discretization choices and convergence
@@ IV Numerical Results
In this section, we present QNM shifts for overtones $n=0,1,2,3$ and multipoles $\ell=2,3,4$ in both even and odd sectors. We work in units $r_{g}=1$ throughout, so that the $n=0$ modes can be compared directly with results of Ref. [1].
@@ IV.1.1 Null operators
It is instructive to compare with the results of Ref. [1] that uses method developed in [32] to compute the QNMs. They have focused on the fundamental modes ($n=0$) with $\ell=2,3,4$ and reported maximal null-operator shifts of order $10^{-4}$ in the even sector and $10^{-6}$ in the odd sector. Over the same set of modes, our EVP implementation reduces these residuals by several orders of magnitude, yielding $|\delta_{\mathrm{EVP}}|\lesssim 10^{-17}$ while simultaneously extending the null test to overtones up to $n=3$ without loss of control. We speculate that the improvement is primarily due to avoiding the series truncation error used In Ref. [1] and  [32]: in EVP we compute the shift directly from the perturbed eigenvalue problem at high precision, and verify convergence so the residual in null cases is driven down to a much lower numerical floor.
@@ IV.1.2 Control Operators
Up to an overall sign (that can be traced to a difference in conventions), the EVP results for $\mathcal{O}_{9}$ and $\mathcal{O}_{10}$ operators gives a fractional shifts $\delta_{\mathrm{EVP}}$ are $\mathcal{O}(10^{0}-10^{2})$. They agree with those of Ref. [1] in magnitude for both the real and imaginary parts upon rounding to the two decimal places reported in their Table I.
@@
Finally, we stress that the EVP computation only probes the linear response of the spectrum, encoded in the first–order coefficients $\omega_{1}$ and hence in $\delta_{\mathrm{EVP}}$. By construction, this method cannot determine the range in which the expansion
$\omega(\zeta)=\omega_{0}+\zeta\,\omega_{1}+\zeta^{2}\omega_{2}+\mathcal{O}(\zeta^{3})$ (47)
is well-approximated by its linear truncation.
@@ IV.2 Leaver method + band reduction
Now we compute the EFT-corrected frequency $\omega(\zeta)$ for some values $\zeta=\alpha\,10^{-p}$ with $p\in\{3,4,\ldots,15\}$ and $\alpha\in\left\{1,\frac{3}{4},\frac{1}{2},\frac{1}{4}\right\}$, keeping only $\zeta\in[10^{-15},10^{-3}]$.
@@
The EFT linear expansion in $\zeta$ is only meaningful as long as the background is not strongly deformed. An out-of-envelope diagnostic is the correction to the horizon radius $r_{H}$. For example, we find that at $d_{5}\zeta=10^{-2}$ the horizon is shifted by about $12\%$ from $r_{g}$, while for $d_{9}\zeta=10^{-2}$ the shift is about $10\%$. [^2] Pushing to $d_{i}\zeta=10^{-1}$ is not even physically meaningful, since $r_{H}$ would formally become negative. To remain safely within the linear regime, in what follows we therefore take $\zeta_{\max}=10^{-4}$ as our default upper limit, and treat $\zeta=10^{-3}$ only as a reference point illustrating the onset of linear perturbation failure [^3] In what follows, Wilson coefficients $d_{i}\in\{0,1\}$ is a switch that selects the $i$th EFT operator, while $\zeta$ is the explicit perturbation parameter..
@@
Although the underlying field equations are linearized in $\zeta$, the numerical procedure used to extract $\omega(\zeta)$ via Leaver’s method can effectively introduce higher-order dependence on $\zeta$. Two immediate sources are: (i) the Gaussian-elimination band-reduction step, which combines $\zeta$-dependent terms and therefore generates products that promote higher powers of $\zeta$  [^4] A strictly linear-in-$\zeta$ reduction can be enforced by truncating to $\mathcal{O}(\zeta)$ after each elimination step, nevertheless, we do not impose this truncation in our implementation., and (ii) the nonlinear Newton solver employed to locate the roots of the continued fraction, which can also imprint nonlinear $\zeta$-dependence through its iterative updates. As a result, the numerically extracted frequency is more accurately viewed as an expansion
@@
where $a_{1}$ is the linear coefficient, while $a^{(1)}_{2},a^{(1)}_{3}$$\ldots$ denote effective higher-order contributions induced by the Leaver extraction procedure, even when the input perturbation is truncated at $\mathcal{O}(\zeta)$. Consequently, estimating $a_{1}$—especially in cases where the linear contribution is suppressed or vanishes for a given operator—is a subtle task.
Diagnostics.
@@
IV.2.1 What is the leading power?
@@
where the coefficient $C\in\mathbb{C}$ are generally complex.
On a chosen coupling window $\zeta_{\min},\zeta_{\max}]$, we then infer the leading scaling via two complementary steps: (i) estimate a continuous effective exponent $p_{\rm eff}$ from log–log slopes of the amplitude $\|\delta_{\mathrm{Leaver}}\|$ across adjacent $\zeta$ values, and we then average these local slope estimates across the $\zeta$ points within the chosen fitting window $\zeta_{\min},\zeta_{\max}]$ to obtain the reported $p_{\rm eff}$ (with scatter across the window used to define $\sigma_{p}$), and (ii) test candidate integers $p\in\{1,2,3\}$ by rescaling the data by $\zeta^{p}$ and selecting the $p_{\star}$ that makes the rescaled complex series as flat as possible. Finally, with $p_{\star}$ fixed, we obtain the corresponding complex coefficient $C$ from a least-squares fit using the flatness ratio $\sigma/\mu$ reported in the tables where $N$, $\mu$ and $\sigma$ is the sample population, mean and standard deviation respectively inside the fit window.
@@
Upon tightening the window to $\zeta_{\max}=10^{-4}$ and $10^{-5}$ (as shown in Tables 21, 22, 23, and 24), the same $p_{\star}=2$ conclusion is reinforced: $p_{\rm eff}$ moves closer to $2$ and $\sigma/\mu$ decreases, consistent with convergence toward the asymptotic quadratic regime. We therefore identify $p=2$ as the leading power for both null operators in both parities
@@
Tightening the window to $\zeta_{\max}=10^{-4}$ and $10^{-5}$ (as shown in Tables 17, 18, 19, and 20) preserves the same $p_{\star}=1$ selection while reducing uncertainties and significantly decreasing $\sigma/\mu$; the fitted complex coefficient $C$ (reported via $\Re(C)\,|\,\Im(C)$) varies only mildly across windows, consistent with a well-resolved linear leading term.
IV.2.2 Polynomial fits
@@
$M_{123}$ $:\quad\Delta\omega(\zeta)\approx c^{(123)}_{1}\,\zeta+c^{(123)}_{2}\,\zeta^{2}+c^{(123)}_{3}\,\zeta^{3},$
where we use $M_{123}$ only for the null-operator cases.
Moreover, although our calculated QNMs have $\sim 60$ digits of precision and we could in principle include work with $M_{1234}$ as well, we won’t to avoid overfitting  [^6] Nevertheless, when stated explicitly, we also fit $M_{1234}$ as a reference model; it is not used to set our reported bounds or conclusions.. This is the same reason why $M_{123}$ is not going to be used for control cases.
@@
IV.3 Cross Check: EVP Vs Leaver
@@ V Discussion and Conclusion
Next-generation gravitational-wave observatories will enable high-SNR ringdown measurements and make black-hole QNMs increasingly sensitive probes of small, high-scale deviations from GR in the dynamical strong-field regime [2, 4, 6, 40]. A controlled way to parameterize such deviations is the gravitational EFT, in which new physics is encoded in higher-curvature operators suppressed by a high scale [1]. Before interpreting tiny EFT-induced QNM shifts as physical predictions, however, the framework and its numerical extraction must be stress-tested: (i) the problem is intrinsically two-parameter, involving both the GW perturbation order $\epsilon$ and the EFT deformation $\zeta$; and (ii) the expected signals can be comparable to the numerical floor, especially for overtones, unless precision and convergence are tightly controlled.
@@
. On a Ricci-flat background, the operators $\mathcal{O}_{5}$ and $\mathcal{O}_{8}$ are removable by local field redefinitions and therefore induce no first-order ($\mathcal{O}(\zeta)$) vacuum QNM shifts in theory [1, 7]. This provides a sharp, operator-level null prediction against which both solvers can be stress-tested.
@@ Appendix B More Details On Numerical Methods
This gives a numerically stable approximation to $R_{0}(\omega)$, and we then solve the QNM condition $F(\omega)=0$ using a two-point complex secant method. The precise numerical tolerances and parameter choices for $N_{\rm tail}$, $\Delta N$, and $p$ are summarized in the III.2.1.